OpenStudy (anonymous):

Medal for help_ A souvenir makers wants to create a scale model of the empire state Building.The ESB is 1,472 feet tall and has a base with dimensions 286 feet by 286 feet If the model is 6 in tall, approximately what are the dimensions of its base in inches? A. 0.2 in by .02in b. 1 ich by 1 inch c. 2.3 in by 2.3 in d. 1.2 in by 1.2 in

OpenStudy (anonymous):

@sammixboo @greenlegodude57

OpenStudy (anonymous):

im back cis my answer

OpenStudy (anonymous):

@AnswerMyQuestions @kirbykirby

OpenStudy (anonymous):

I don't understand the question.

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@ARMYRANGERS

OpenStudy (armyrangers):

The answer is D

sammixboo (sammixboo):

Well here is what I did to get the answer, it might not be the way you do it, but it is the way I do it, lol. I took 1,472ft and 186ft made them a ratio: \(\frac{ 286 }{ 1,472 }\) : and I also put 6 in ratio form, but with a missing part: \(\frac{n}{6}\) : 1,472 and 6 are at the bottom of the ratio because they are the length of the ESB and 286 is the base. \(\frac{286}{1,472}\) = \(\frac{n}{6}\) \(\frac{286}{1,472}\) = \(\frac{n}{6}\) Now I did \(286*6\) which is \(1,716\) So now I know that \(1,472*n\) has to equal \(1,716\) Now what I did was take \(1,472\) and divide it by \(1,716\) and got an answered rounded to \(1in\) Idk if you will understand it or not, but I got B. 1 in. by 1 in. I could be wrong though.

OpenStudy (armyrangers):

Sam the bigger # is for the hight they are asking for the base and the height has to be six inches

OpenStudy (armyrangers):

If you do 1472/245 you get 6.008 so then you take 286/245 and you get 1.16 so you round up any you 1.2 by 1.2

OpenStudy (armyrangers):

@sammixboo

sammixboo (sammixboo):

Okie Dokie